OpenStudy (anonymous):

log(x-5)=log4-log(x+2)

OpenStudy (btaylor):

How do you condense the right side so that you can eliminate the log?

OpenStudy (btaylor):

let me give you a hint: \[\log(a)-\log(b)=\log \left( \frac{a}{b} \right)\]

OpenStudy (solomonzelman):

\[\huge\color{blue}{ \log(x-5)=\log(4)-\log(x+2) } \]\[\huge\color{blue}{ \log(x-5)=\log(\color{red} { 4 })-\log(\color{red} { x+2 }) } \]\[\huge\color{blue}{ \log(x-5)=\log(\color{blue} { \frac{\color{red} { 4 }}{\color{red} { x+2 }} }) } \] \[\huge\color{blue}{ x-5=\frac{4}{x+2} } \] \[\huge\color{blue}{ \frac{x-5}{1}=\frac{4}{x+2} } \] \[\huge\color{blue}{ (x-5)(x+2)=1 \times 4 } \] \[\huge\color{blue}{x^2-3x-10=4 } \] can you solve this quadratic?